A chemist needs 18L of a 20% solution of alcohol. she has on hand a 15% solution as well as a 30% solution. How many L of the 15% solution and the 30% solution should she mix to get the 20% solution needed?
Answer by checkley77(12844)

.30x+.15(18-x)=.20*18
.30x+2.7-.15x=3.6
.15x=3.6-2.7
.15x=.9
x=.9/.15
x=6 L. of 30% solution.
Proof:
.30*6+.15(18-6)=.20*18
1.8+.15*12=3.6
1.5+1.8=3.6
3.6=3.6